Corrections to An introduction to Kato’s Euler systems

In the final theorem 5.2.7, a crucial statement was omitted: the constant C_{E}^{±} belongs to Q^{∗}. (It is a non-zero rational multiple of the constant C′ in the previous equation.)

Last modified 16 April 1999